Refunding Analysis

Mullet Technologies is considering whether or not to refund a $50 million, 15% coupon, 30-year bond issue that was sold 5 years ago. It is amortizing $3 million of flotation costs on the 15% bonds over the issue's 30-year life. Mullet's investment banks have indicated that the company could sell a new 25-year issue at an interest rate of 9% in today's market. Neither they nor Mullet's management anticipate that interest rates will fall below 9% any time soon, but there is a chance that rates will increase.

A call premium of 15% would be required to retire the old bonds, and flotation costs on the new issue would amount to $6 million. Mullet's marginal federal-plus-state tax rate is 40%. The new bonds would be issued 1 month before the old bonds are called, with the proceeds being invested in short-term government securities returning 7% annually during the interim period.

  1. Conduct a complete bond refunding analysis. What is the bond refunding's NPV? Do not round intermediate calculations. Round your answer to the nearest cent.

    $  

  2. What factors would influence Mullet's decision to refund now rather than later?
Answer #1
Current bond issue information
Par value 50,000,000.00
coupon rate 15%
original maturity 30
remaining maturity                           25
original flotation costs 3,000,000.00
Call premium 15%
Tax rate 40%
New issue information
Coupon rate 9%
maturity                           25
flotation costs 6,000,000.00
Time between issues (months) 1
rate on surplus funds (annual) 7%
a. Conduct a complete bond refunding analysis. What is the bond refunding's NPV?  
Initial investment outlay to refund old issue:
Call premium on old issue = 50,000,000*15% = 7,500,000.00
After-tax call premium = 7,500,000*(1-40%) = A 4,500,000.00
New flotation cost = B 6,000,000.00
Old flotation costs already expensed = (3,000,000/30)*(30-25) = 500,000.00
Remaining flotation costs to expense = 3,000,000-500,000 = 2,500,000.00
Tax savings from old flotation costs = 40%*2,500,000 = C 1,000,000.00 You get to expense the remaining flotation costs
Additional interest on old issue after tax = 50,000,000*15%*(1/12)*(1-40%) = D 375,000.00 This is interest paid on the old bond issue between when the new bonds are issued and the old bonds are retired
Interest earned on investment in T-bonds after tax = (50,000,000*7%*1/12)*(1-40%) = E 175,000.00 This is interest earned on the proceeds from the new bonds before they are used to pay off the old bonds.
Total investment outlay = A+B-C+D-E = 9,700,000.00
Annual Flotation Cost Tax Effects:
Annual tax savings on new flotation = 6,000,000*40%/25 = F 96,000.00
Tax savings lost on old flotation = 3,000,000*40%/30 G 40,000.00
Total amortization tax effects = F-G = H 56,000.00
Annual interest savings due to refunding:
Annual after tax interest on old bond = 50,000,000*15%*(1-40%) I 4,500,000.00
Annual after tax interest on new bond = 50,000,000*9%*(1-40%) J 2,700,000.00
Net after tax interest savings = I - J = K 1,800,000.00
Annual cash flows = H+K = 1,856,000.00
NPV of bond refunding decision = -9,700,000-1,856,000*(((1-1.07^-25))/0.07) 11,929,050.38
b. What factors would influence Mullet's decision to refund now rather than later?
Mullet will do the refund transaction now because it has view that interest rate will not fall below 9% rather it may increase in nearby future.
So, It is beneficial to do the refund transaction now as if interest goes over 12.41% it will give negative NPV as 12.41% is break-even interest rate.
and if interest rate remains between 9% to 12.41% it will give positive returns to Mullet.